OpenStudy (anonymous):

What is the domain of the following function? f(x) = √-x - 1- 3 f(x) ≤ -1 x ≤ 1 f(x) ≤ 1 x ≤ -1

OpenStudy (anonymous):

which part is in root

OpenStudy (anonymous):

all of it. -x-1-3 i s in root

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ f(x)=\sqrt{-x-1}-3}\)

OpenStudy (solomonzelman):

The domain can be any real number that is less than or equal to -1, so that you don't get an imaginary in the square root.

OpenStudy (anonymous):

she says the whole thing is in root

OpenStudy (solomonzelman):

It can not be, it is silly to say \(\normalsize\color{blue}{ f(x)=\sqrt{-x-1-3}}\)

OpenStudy (solomonzelman):

It should be, \(\normalsize\color{blue}{ f(x)=\sqrt{-x-1}-3}\)

OpenStudy (solomonzelman):

And for that, domain ≤ -1

OpenStudy (anonymous):

*shrugs*

OpenStudy (haseeb96):

so f(x) =sqrt -x -1-3 f(x) = sqrt -x -4 and f(x) =-1 okay left hand side is same so compare right hand side -1=sqrt -x -4 now square on both sides 1 = -x -4 x=-4-1 x=-5 is yur domain of this function and again do it with f(x)=1 okay ?

OpenStudy (solomonzelman):

I want to hear from the asker what the question is exactly like.

OpenStudy (anonymous):

that is sooooooo soooo soooooooo wrong @Haseeb96

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ f(x)=\sqrt{-x-1}-3}\) Domain ≤ -1 \(\normalsize\color{blue}{ f(x)=\sqrt{-x-1-3}}\) Domain ≤ -4

OpenStudy (solomonzelman):

By the way guys, 1) Click and hold ALT 2) click the number code (using the numbers that are on the right of the keyboard, and `NOT` the ones below `F1`, `F2`, `F3`, etc., ) 3) release the ALT number code result ` 2 5 1 ` √ `0 2 1 5 ` × `2 4 6 ` ÷ `7 5 4 ` ≥ `7 5 5` ≤ `2 4 1 ` or `7 5 3` ± `2 4 7` ≈ `0 1 8 5 ` ¹ `2 5 3 ` ² `0 1 7 9 ` ³

OpenStudy (anonymous):

®

OpenStudy (anonymous):

omg really i look away for a couple mins and i come back to this? lol its -1. end of discussion. thank you everyone.

OpenStudy (anonymous):

hes kinda a retrice

OpenStudy (anonymous):

haha retrice

OpenStudy (anonymous):

haha it changes butt cuss word into retrace.. lol

OpenStudy (solomonzelman):

No it is -1, or anything that is less than -1. Assuming that you meant \(\normalsize\color{blue}{ f(x)=\sqrt{-x-1}-3}\) initially.